Find an expression for a square matrix A satisfying A squared

= In, where In is the n × n identity matrix. Give 3

examples for the case n = 3. 


1
Expert's answer
2022-06-01T03:33:07-0400

Matrix "A" is an involutory matrix, if "A^2=I_n" .

There is no formula for this type matrix, but there many examples.

All matrix "A=\begin{pmatrix}
\pm 1 &0&0&…&0
\\
0& \pm1&0&…&0
\\
0&0&\pm1&…&0
\\
\vdots &\vdots &\vdots& \ddots &0
\\
0&0&0&0&\pm1
\end{pmatrix}" satisfy "A^2=I_n" .


Examples of "3\times 3" matrices:

"A=\begin{pmatrix}
1&0&0\\0&1&0\\0&0&1
\end{pmatrix}"

"B=\begin{pmatrix}
-1&0&0\\0&1&0\\0&0&-1
\end{pmatrix}"

"C=\begin{pmatrix}
1&0&0\\0&0&-1\\0&-1&0
\end{pmatrix}"